OpenStudy (anonymous):

THE last one please help!! P(x) = x4 - 7x3 + 6x2. Describe the graph by discussing the rising, falling, zero points and the general shape of the graph.

OpenStudy (anonymous):

what level of math?

OpenStudy (anonymous):

if its lower lvl, find the zeroes by facctoring it out. x^2(x^2-7x+6) x^2=0,x=0 (x^2-7x+6)=(x-1)(x-6)=0, x=1, x=6

OpenStudy (anonymous):

if in calc, take the derivitive find the zeroes of that and that will give you the point wher it changes from rising to falling.